# Distillation versus Contrastive Learning: How to Train Your Rerankers

### v2

@@ 4.1 Main Results
BM25 Lin et al. (2021) - - $|\mathcal{D}|$ 18.4 50.6 48.0
CoCondenser Gao and Callan (2022) 110M - $|\mathcal{D}|$ 38.2 71.7 68.4
RepLlama2 Ma et al. (2023) 7B - $|\mathcal{D}|$ 41.2 74.3 72.1
CSPLADE Xu et al. (2025a) 8B - $|\mathcal{D}|$ 41.3 74.1 72.8
@@
cross-SimLM Wang et al. (2022a) 110M bi-SimLM 200 43.7 74.6 72.7
RankT5 Zhuang et al. (2023) 220M GTR 1000 43.4 - -
RankMamba Xu et al. (2025c) 1.3B BGE 100 38.6 75.8 74.0
@@
[…]
Out-of-domain results.
We report the out-of-domain evaluation results in Table˜2. The baseline methods’ results are deferred to Appx.˜C.
Among all models with KD training strategy, RankQwen-3B-KD achieves strong out-of-domain performance, averaging 57.5 NDCG@10 over 13 BEIR datasets, improving over RankQwen-7B-CL teacher. However, we notice that RankQwen-7B-CL’s underperformance is mainly due to Quora duplicate question retrieval — a symmetric retrieval task different from the asymmetric MS MARCO trainset. The performance is 55.5 versus 56.2 excluding Quora, suggesting that the student model still cannot improve over the teacher Gou et al. (2021); Gholami and Omar (2023).
An important observation is that RankQwen-7B-KD leads to performance degradation in out-of-domain evaluation (55.8 vs RankQwen-3B-KD’s 57.6), while their performances on in-domain MS MARCO datasets are comparable. Prior studies have also noted that knowledge distillation suffers from overfitting and poor out-of-domain generalization Gou et al. (2021); Yuan et al. (2020); Yun et al. (2020). Our experimental results suggest that when distilling from a larger, more capable teacher models (RankQwen-7B-CL in our case), the student still achieve performance improvement compared to contrastive training from scratch.
Scaling model sizes.
Scaling has been proven effective in IR tasks (Neelakantan et al., 2022; Muennighoff et al., 2024b; Xu et al., 2025b; Zhu et al., 2023, inter alia). We also compare the model size scaling trend between CL versus KD for passage reranking in Fig.˜1. We make two observations:
  1. 1.
    We observe that contrastive learning demonstrates a clear scaling trend, i.e., performance improves with increased model size when trained on the same data, as also reported by prior works Ma et al. (2023); Neelakantan et al. (2022); Muennighoff (2022); Zhuang et al. (2024a).
  2. 2.
Retriever Teacher Contrastive Learning Knowledge Distillation
Dataset 110M 7B 0.5B 3B 0.5B 3B
[…]
@@ 4.2 Scaling Training Data
In §˜4.1, we focus on a controlled experiment with various model sizes. Now we examine the efficacy of knowledge distillation when scaling the axis of training data. Notice that doing so leads to the decoupling of the retriever and reranker as we can no longer mine hard negatives from the retriever on MS MARCO passage trainset. We use a recently released trainset, RLHN Thakur et al. (2025), which consists of 649K $(q,d^{+},D^{-})$ training samples pruned from the larger BGE trainset Xiao et al. (2023). Specifically, RLHN comprises train samples from the following dataset: MS MARCO passage Bajaj et al. (2016), Arguana Wachsmuth et al. (2018), FEVER Thorne et al. (2018), FiQA Maia et al. (2018), SCIDOCS Cohan et al. (2020), HotpotQA Yang et al. (2018) and NQ Kwiatkowski et al. (2019). In this case, BEIR benchmark as a whole is no longer considered out-of-domain evaluation. We use a lightweight intfloat/e5-base-v2 retriever Wang et al. (2022b) pretrained on unlabeled text pairs, then finetuned with the combined MS MARCO passage and NQ training mixture.
We use the similar training strategy as §˜3 (the only difference is the ranker training data): we first train 7B-scale reranker with Qwen2.5-7B and contrastive learning, then train smaller rerankers with knowledge distillation, with the reranking scores on RLHN dataset labeled by the teacher, as well as the contrastive learning counterparts. We train 0.5B and 3B rerankers, as we notice Qwen2.5-1.5B fail to converge in the contrastive learning setting. We name the trained rerankers as RankQwen-{0.5B, 3B, 7B}-CL-RLHN and RankQwen-{0.5B, 3B}-KD-RLHN, respectively.
@@
In  Fig.˜2 we further analyze BEIR in-domain (7 datasets used in training) and out-of-domain (7 unseen datasets) performance with varying training data sizes. As the amount of training data increases, both CL and KD show improved in-domain performance, with CL eventually outperforming KD. In contrast, KD consistently outperforms CL on the OOD datasets, although the benefits of increasing training data are diminishing. We leave a more in-depth investigation of the in-domain versus OOD performance gap to future work.
5 Related Works
In this paper, we focus exclusively on the architecture of cross-encoder pointwise reranker Nogueira and Cho (2019); Nogueira et al. (2020); Zhuang et al. (2024b). The reranker model processes a query and a document simultaneously, allowing the self-attention mechanism to explicitly model the interactions between their tokens throughout the layers. Nogueira and Cho (2019) first demonstrated the efficacy of this approach by fine-tuning BERT as a text pair classifier. The fine-tuning of the BERT-type model as cross-encoder reranker has later been extended to encoder-decoder architectures, such as T5 Raffel et al. (2020); Nogueira et al. (2020), and decoder-only models like Llama Touvron et al. (2023); Ma et al. (2023).
The training objective for a cross-encoder is to accurately discriminate between relevant and non-relevant documents for a given query. In practice, the non-relevant documents are usually hard negatives mined from first-stage retrievers (Nogueira et al., 2019; Gao et al., 2021; Boytsov et al., 2022, inter alia). Two primary training strategies are widely used to train cross-encoder reranker. The first is contrastive learning, which trains the model directly on ground-truth labels to maximize the distinction between positive and negative examples Gao et al. (2021); Zhuang et al. (2023). The second strategy is knowledge distillation, which involves training a student reranker to mimic the behavior and performance of a larger, more capable teacher model Hofstätter et al. (2020); Baldelli et al. (2024); Schlatt et al. (2025). Knowledge distillation (KD, Buciluǎ et al., 2006; Hinton et al., 2015) facilitates the transfer of knowledge from a large, complex “teacher” model to a smaller, more efficient “student” model. The goal is to enable the student to mimic the teacher’s output, thereby inheriting its predictive capabilities at a reduced computational cost. Within the context of training neural IR models, knowledge distillation has been effectively applied to enhance both bi-encoder retriever (Hofstätter et al., 2020; Formal et al., 2021, inter alia) and cross-encoder reranker Baldelli et al. (2024); Schlatt et al. (2025). In this work, we aim to compare these two training strategies in a controlled experimental setting to evaluate their strengths and weaknesses.
@@ 6 Conclusion and Future Work
In this paper, we compared the effectiveness of two training strategies, i.e., contrastive learning and knowledge distillation in the context of training text rerankers. With a rigorously controlled experimental setup, we find that when distilling from a larger, more capable teacher model, rerankers trained with knowledge distillation achieve better in-domain and out-of-domain reranking performances compared to contrastive training strategy, and the observation is consistent across different model scales and language model architectures. Our future work will investigate a more optimized way to combine two training strategies, as well as to improve the robustness and out-of-domain generalization of knowledge distillation training.
Limitations
Given the limited computational resources, we are unable to scale the reranker training to >10B models such as Qwen2.5-14B. The observation made in this work may be subject to change for stronger base models. How to improve robustness and out-of-domain generalization of the knowledge distillation training has been extensively investigated by prior works in other NLP domains Utama et al. (2020); Stacey and Rei (2024); Wang et al. (2023), we leave investigation for reranking to future works.
This paper focuses on empirical experiments on public benchmarks. We believe this paper do not incur potential risks.
@@ Appendix A Dataset Artifacts and Licenses
Four of the datasets we used in experiments (NFCorpus Boteva et al. (2016), FiQA-2018 Maia et al. (2018), Quora [^2] https://www.kaggle.com/c/quora-question-pairs, Climate-Fever Diggelmann et al. (2020)) do not report the dataset license in the paper or a repository. For the rest of the datasets, we list their licenses below:
  • MS MARCO Bajaj et al. (2016): MIT License for non-commercial research purposes.
  • ArguAna Wachsmuth et al. (2018): CC BY 4.0 license.
  • DBPedia Hasibi et al. (2017): CC BY-SA 3.0 license.
  • FEVER Thorne et al. (2018): CC BY-SA 3.0 license.
  • HotpotQA Yang et al. (2018): CC BY-SA 4.0 license.
  • NQ Kwiatkowski et al. (2019): CC BY-SA 3.0 license.
  • SCIDOCS Cohan et al. (2020): GNU General Public License v3.0 license.
  • SciFact Wadden et al. (2020): CC BY-NC 2.0 license.
  • TREC-COVID Voorhees et al. (2021): "Dataset License Agreement".
  • Touche-2020 Bondarenko et al. (2020): CC BY 4.0 license.
Appendix B Hyperparameters
For all our training runs, we use a similar set of optimized hyperparameters identified from prior works and our preliminary experiments Ma et al. (2023); Xu et al. (2025a), only ablating learning rate to reduce the effect of overfitting. We use LoRA rank=16 and $\alpha$=32. We use AdamW optimizer, learning rate ranging from 3e-5 to 1e-4 with linear warmup and cool down. As we train all models on the train dataset for 1 epoch, we find learning rate is the most important hyperparameter to control overfitting. We use 8 GPUs with per device batch 4, gradient accumulation steps 4, which leads to a global batch size of 128. For each $(q_{i},d_{i}^{+})$ pair, we use $k=15$ negatives as we find increasing to 313131 negatives may lead to instable contrastive learning. For RepQwen training, we use in-batch negatives; while for reranker training, we focus solely on each $(q_{i},d_{i}^{+})$ pair’s own hard negatives.
Appendix C Baseline Results

### v3

@@ 5.1 Main Results
Table 1: Results for passage reranking in-domain evaluation. We mark best results in each section bold; $\dagger$} indicates the overall best result and $\ddagger$} indicates the best result among our trained models; for models in Retrieval and Reranking baseline sections, denotes the model we trained. Note RankQwen-7B-CL is used as the teacher for distillation.
@@
BM25 lin2021pyserini - - $|\mathcal{D}|$ 18.4 50.6 48.0
CoCondenser gao2022unsupervised 110M - $|\mathcal{D}|$ 38.2 71.7 68.4
RepLlama2 ma2023fine 7B - $|\mathcal{D}|$ 41.2 74.3 72.1
CSPLADE xu2025csplade 8B - $|\mathcal{D}|$ 41.3 74.1 72.8
@@
cross-SimLM wang2022simlm 110M bi-SimLM 200 43.7 74.6 72.7
RankT5 zhuang2023rankt5 220M GTR 1000 43.4 - -
RankMamba xu-etal-2025-statespacemodelsarestrongtextrerankers 1.3B BGE 100 38.6 75.8 74.0
@@
Table 2: Results for passage reranking out-of-domain evaluation. 0.5B, 1.5B, 3B, 7B 2B correspond to RankQwen-0.5B, RankQwen-1.5B, RankQwen-3B, RankQwen-7B and RankRGemma-2B, respectively. We also report results without Quora dataset as Quora duplicate question detection is a symmetric retrieval task, not aligning with the asymmetric web search task of the training set. We mark best results in each section bold; $\dagger$} indicates the overall best result including the teacher model, $\ddagger$} indicates the best result excluding teacher.
@@
[…]
FiQA 45.9 37.9 42.2 43.2 44.1 34.7 45.4 $6{\dagger}{\ddagger} 4$4.2 44.7
HotpotQA 76.1 72.5 74.8 76.2 74.4 74.4 $6.7{\dagger}{\ddagger} 72$.5 76.6
NFCorpus 31.5 33.0 29.3 25.9 $4{\dagger}{\ddagger} 3$2.2 32.3 33.3 32.3 34.2
[…]
SCIDOCS 19.1 16.9 16.2 17.3 18.4 17.8 ${\dagger} 19.2{\dagger}{\ddagger}$ 19.1 18.6
[…]
TREC-COVID 85.0 81.9 86.3 86.8 85.1 80.7 85.6 86.3 $87.4{\dagger}{\ddagger}$
[…]
Average 56.5 53.3 55.3 56.2 55.2 53.1 $7.6{\dagger}{\ddagger} 55$.8 56.5
[…]
Out-of-domain results.
We report the out-of-domain evaluation results in Table˜2. The baseline methods’ results are deferred to Appendix˜C.
Among all models with KD training strategy, RankQwen-3B-KD achieves strong out-of-domain performance, averaging 57.5 NDCG@10 over 13 BEIR datasets, improving over RankQwen-7B-CL teacher. However, we notice that RankQwen-7B-CL’s underperformance is mainly due to Quora duplicate question retrieval — a symmetric retrieval task different from the asymmetric MS MARCO trainset. The performance is 55.5 versus 56.2 excluding Quora, suggesting that the student model still cannot improve over the teacher gou2021knowledgedistillationsurvey; Gholami2023canastudentllmperformaswellasteacher.
An important observation is that RankQwen-7B-KD leads to performance degradation in out-of-domain evaluation (55.8 vs RankQwen-3B-KD’s 57.6), while their performances on in-domain MS MARCO datasets are comparable. Prior studies have also noted that knowledge distillation suffers from overfitting and poor out-of-domain generalization gou2021knowledgedistillationsurvey; Yuan2020Revisitingknowledgedistillation; Yun2020Regularizingclasswisepredictionsviaselfknowledgedistillation. Our experimental results suggest that when distilling from a larger, more capable teacher models (RankQwen-7B-CL in our case), the student still achieve performance improvement compared to contrastive training from scratch.
Scaling model sizes.
Scaling has been proven effective in IR tasks (neelakantan2022text; muennighoff2022sgpt; xu2025surveyofmodelarchitectures; zhu2023llmforirsurvey, inter alia). We also compare the model size scaling trend between CL versus KD for passage reranking in Figure˜1. We make two observations:
  1. 1.
    We observe that contrastive learning demonstrates a clear scaling trend, i.e., performance improves with increased model size when trained on the same data, as also reported by prior works ma2023fine; neelakantan2022text; muennighoff2022sgpt; zhuang-etal-2024-promptreps; xu2024rankmamba; muennighoff2024gritlm.
  2. 2.
Table 3: Results with on BEIR benchmark (including MS MARCO passage Dev) with RLHN training mixture. We report NDCG@10 as the performance metric. We mark best result in each row bold; $\dagger$} indicates the overall best result excluding the 7B-sized teacher model.
Retriever Teacher Contrastive Learning Knowledge Distillation
Dataset 110M 7B 0.5B 3B 0.5B 3B
[…]
Arguana 44.5 79.1 65.2 $7{\dagger} 6$3.7 75.6
@@ 5.2 Scaling Training Data
In §˜5.1, we focus on a controlled experiment with various model sizes. Now we examine the efficacy of knowledge distillation when scaling the axis of training data. Notice that doing so leads to the decoupling of the retriever and reranker as we can no longer mine hard negatives from the retriever on MS MARCO passage trainset. We use a recently released trainset, RLHN thakur2025fixingdatathathurtsperformance, which consists of 649K $(q,d^{+},D^{-})$ training samples pruned from the larger BGE trainset bge_embedding. Specifically, RLHN comprises train samples from the following dataset: MS MARCO passage bajaj2016ms, Arguana wachsmuth2018retrieval, FEVER thorne2018fever, FiQA maia201818, SCIDOCS cohan2020specter, HotpotQA yang2018hotpotqa and NQ kwiatkowski-etal-2019-natural. In this case, BEIR benchmark as a whole is no longer considered out-of-domain evaluation. We use a lightweight intfloat/e5-base-v2 retriever wang2022e5embedding pretrained on unlabeled text pairs, then finetuned with the combined MS MARCO passage and NQ training mixture.
We use the similar training strategy as §˜4 (the only difference is the ranker training data): we first train 7B-scale reranker with Qwen2.5-7B and contrastive learning, then train smaller rerankers with knowledge distillation, with the reranking scores on RLHN dataset labeled by the teacher, as well as the contrastive learning counterparts. We train 0.5B and 3B rerankers, as we notice Qwen2.5-1.5B fail to converge in the contrastive learning setting. We name the trained rerankers as RankQwen-{0.5B, 3B, 7B}-CL-RLHN and RankQwen-{0.5B, 3B}-KD-RLHN, respectively.
@@
In  Figure˜2 we further analyze BEIR in-domain (7 datasets used in training) and out-of-domain (7 unseen datasets) performance with varying training data sizes. As the amount of training data increases, both CL and KD show improved in-domain performance, with CL eventually outperforming KD. In contrast, KD consistently outperforms CL on the OOD datasets, although the benefits of increasing training data are diminishing. We leave a more in-depth investigation of the in-domain versus OOD performance gap to future work.
@@ 6 Conclusion and Future Work
In this paper, we compared the effectiveness of two training strategies, i.e., contrastive learning and knowledge distillation in the context of training text rerankers. With a rigorously controlled experimental setup, we find that when distilling from a larger, more capable teacher model, rerankers trained with knowledge distillation achieve better in-domain and out-of-domain reranking performances compared to contrastive training strategy, and the observation is consistent across different model scales and language model architectures. Along this direction, our future work will investigate a more optimized way to combine two training strategies, and to improve the robustness and out-of-domain generalization of knowledge distillation training.
Recent works have highlighted the efficacy of a new knowledge distillation paradigm — on-policy distillation agarwal2024policy; yang2025qwen3technicalreport; lu2025onpolicydistillation. Mapping to the classical information retrieval tasks like reranking, we note that existing training strategies could be generally considered off-policy from the reinforcement learning perspective, where the model is optimized to imitate outputs from external source, either from click information joachims2017accurately; tu2025reinforcementlearningrankusing, human annotation bajaj2016ms or LLMs zhang2025qwen3embedding, which may encounter the train-inference distribution mismatch problem described by agarwal2024policy, and can be traced back to extensive RL literature. We believe a promising future direction is to study how reranking can be improved by on-policy training, where the model gets rewarded based on its current policy.
Limitations and Potential Risks
Given the limited computational resources, we are unable to scale the reranker training to >10B models such as Qwen2.5-14B. The observation made in this work may be subject to change for stronger base models. How to improve robustness and out-of-domain generalization of the knowledge distillation training has been extensively investigated by prior works in other NLP domains Utama2020Mindthetradeoff; stacey-rei-2024-distilling; Wang2023Serialcontrastiveknowledgedistillation, we leave investigation for reranking to future works. This paper focuses on empirical experiments on public benchmarks. We believe this paper do not incur potential risks.
@@ Appendix A Dataset Artifacts and Licenses
Four of the datasets we used in experiments (NFCorpus boteva2016full, FiQA-2018 maia201818, Quora [^4] https://www.kaggle.com/c/quora-question-pairs, Climate-Fever diggelmann2020climate) do not report the dataset license in the paper or a repository. For the rest of the datasets, we list their licenses below:
  • MS MARCO bajaj2016ms: MIT License for non-commercial research purposes.
  • ArguAna wachsmuth2018retrieval: CC BY 4.0 license.
  • DBPedia hasibi2017dbpedia: CC BY-SA 3.0 license.
  • FEVER thorne2018fever: CC BY-SA 3.0 license.
  • HotpotQA yang2018hotpotqa: CC BY-SA 4.0 license.
  • NQ kwiatkowski-etal-2019-natural: CC BY-SA 3.0 license.
  • SCIDOCS cohan2020specter: GNU General Public License v3.0 license.
  • SciFact wadden2020fact: CC BY-NC 2.0 license.
  • TREC-COVID voorhees2021trec: "Dataset License Agreement".
  • Touche-2020 bondarenko2020overview: CC BY 4.0 license.
@@
Appendix B Hyperparameters
For all our training runs, we use a similar set of optimized hyperparameters identified from prior works and our preliminary experiments ma2023fine; xu2025csplade, only ablating learning rate to reduce the effect of overfitting. We use LoRA rank=16 and $\alpha$=32. We use AdamW optimizer, learning rate ranging from 3e-5 to 1e-4 with linear warmup and cool down. As we train all models on the train dataset for 1 epoch, we find learning rate is the most important hyperparameter to control overfitting. We use 8 GPUs with per device batch 4, gradient accumulation steps 4, which leads to a global batch size of 128. For each $(q_{i},d_{i}^{+})$ pair, we use $k=15$ negatives as we find increasing to $31$ negatives may lead to instable contrastive learning. For RepQwen training, we use in-batch negatives; while for reranker training, we focus solely on each $(q_{i},d_{i}^{+})$ pair’s own hard negatives.
Appendix C Baseline Results